### v1

# RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation

Abstract
Code generation has attracted increasing attention with the rise of Large Language Models (LLMs). Many studies have developed powerful code LLMs by synthesizing code-related instruction data and applying supervised fine-tuning. However, these methods are limited by teacher model distillation and ignore the potential of iterative refinement by self-generated code. In this paper, we propose Adaptive Critique Refinement (ACR), which enables the model to refine itself by self-generated code and external critique, rather than directly imitating the code responses of the teacher model. Concretely, ACR includes a composite scoring system with LLM-as-a-Judge to evaluate the quality of code responses and a selective critique strategy with LLM-as-a-Critic to critique self-generated low-quality code responses. We develop the RefineCoder series by iteratively applying ACR, achieving continuous performance improvement on multiple code generation benchmarks. Compared to the baselines of the same size, our proposed RefineCoder series can achieve comparable or even superior performance using less data.
RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation
Changzhi Zhou1thanks:   Work done during internship at Meituan. Equal contribution.,   Xinyu Zhang [^1] footnotemark: 1,   Dandan Song1thanks:   Corresponding Author.,  Xiancai Chen2, Wanli Gu3, Huipeng Ma1, Yuhang Tian1, Mengdi Zhang3, Linmei Hu122footnotemark: 2 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China 2Peking University   3Meituan zhouchangzhi97@gmail.com   {sdd,hulinmei}@bit.edu.cn
1 Introduction
Code generation, also called program synthesis, is a key application area of Large Language Models (LLMs) and has attracted significant attention from the research community Gulwani et al. (2017); Chen et al. (2021). Numerous studies focus on developing code-specific pre-trained models, such as CodeLlama Roziere et al. (2023), DeepSeek-Coder Guo et al. (2024), and Qwen-Coder Hui et al. (2024). These LLMs demonstrate strong capabilities in code understanding owing to their pre-training on large-scale code corpus. However, they must undergo post-training to become user-friendly and effective for code-related tasks.
Figure 1: Comparison of two fine-tuning paradigms. DFT uses code distilled by teacher model, whereas ACR adaptively constructs distinct data formats infusing self-generated code and external critique.
As a popular post-training technology, Distillation-based Fine-Tuning (DFT) leverages code instruction data synthesized by powerful teacher models (often proprietary LLMs like GPT-4) to fine-tune code LLMs, as shown in Figure 1. Many works Zhang et al. (2024b) based on this paradigm have been proposed recently. For example, Code Alpaca Chaudhary (2023) initially establishes this synthesis paradigm using Self-Instruct Wang et al. (2023). Subsequently, Code-related Evol-Instruct Luo et al. (2024) prompts LLMs to evolve more intricate code instructions, while OSS-Instruct Wei et al. (2024b) generates more realistic code instructions based on open-source code snippets from The Stack Kocetkov et al. (2022). In contrast, OpenCodeInterpreter Zheng et al. (2024a) enables code LLMs to learn from compiler diagnostics and human feedback by constructing multi-turn feedback data. These diverse synthesis strategies have significantly advanced the development of code LLMs. However, this paradigm of teacher model distillation inevitably suffers from a critical limitation: the performance of the student model largely relies on the teacher model, ignoring the potential of iterative refinement by self-generated code.
Apart from the research on code instruction fine-tuning, some studies Chen et al. (2023); Huang et al. (2023a); Chen et al. (2024, 2025) have explored the ability of LLMs to refine self-generated code during inference. For example, Self-Debugging Chen et al. (2024) iteratively debugs self-generated code using feedback from the code executor. However, these methods that freeze parameters are essentially a form of prompt engineering with external feedback. They cannot improve the intrinsic one-off code generation ability of LLMs. Besides, multiple calls to LLMs increase the inference latency.
In this paper, we propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm to enhance the intrinsic one-shot code generation ability of LLMs, and on the basis of which we develop a series of code LLMs named RefineCoder. ACR goes beyond traditional teacher distillation by using both self-generated code and external critique to refine the model. Specifically, ACR tailors the distinct new fine-tuning dataset by first scoring and critiquing the self-generated multiple code responses, and then adaptively forms new samples according to the scoring and critiquing results. The two types of samples are shown in Figure 1. Based on the new dataset, we fine-tune RefineCoder models. Finally, we iteratively applying the above ACR process to achieve continuous improvement in code generation capabilities of RefineCoder. Notably, in ACR, we design a composite scoring system and a selective critiquing strategy to better score code responses and selectively critique low-quality code. The former combines LLM-as-a-Judge, an Elo rating mechanism Elo and Sloan (1978), and a code executor for accurate and comprehensive evaluation. The latter adaptively constructs two types of data by comparing the scores of self-generated code with those of original code. When self-generated code performs worse than the original code, an LLM-as-a-Critic is introduced to provide specific critiques.
We conduct iterative ACR based on DS-Coder-6.7B-Base Guo et al. (2024) and Qwen2.5-Coder-7B-Base Hui et al. (2024), resulting in the RefineCoder series. After three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve average pass@1 improvements of 2.4 and 3.0, respectively, on the HumanEval (+), MBPP (+), LiveCodeBench, and the BigCodeBench-hard benchmark Chen et al. (2021); Austin et al. (2021); Liu et al. (2023); Jain et al. (2024); Zhuo et al. (2024). The key contributions of this paper are as follows:
1) We propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm that refines code LLMs with self-generated code and external critique, on the basis of which we develop a series of strong code LLMs (RefineCoder).
2) To ensure the efficacy of ACR, we design a composite scoring system with LLM-as-a-Judge and a selective critique strategy with LLM-as-a-Critic to score and critique self-generated code.
3) Experimental results from the RefineCoder series show that iterative ACR continuously enhances code generation performance. After three iterations, the RefineCoder series achieves comparable or even superior performance than baselines of the same size while requiring less data.
2 Related Work
LLMs for Code Generation LLMs have shown exceptional code understanding abilities due to extensive pre-training on code-related corpora. Numerous models like GPT-4o OpenAI (2024a), Gemini Google (2024), Claude-3.5 Anthropic (2024), Qwen2.5-Coder Hui et al. (2024), and DeepSeek-Coder Guo et al. (2024) exhibit strong performance on code generation benchmarks. Recently, the release of OpenAI o1 OpenAI (2024b) and DeepSeek-R1 DeepSeek-AI (2025) has spurred a surge in research on deep reasoning LLMs, achieving expert-level performance on competitive programming problems (e.g., CodeForces) and further advancing LLMs in the code domain.
Figure 2: Overview of ACR in the $t$-th iteration. (1) Sampling: The model $M_{t}$ samples $n$ code responses $\{y^{i}\}_{i=1}^{n}$ with high temperature for a programming instruction $x$ in the dataset $D_{t}$. (2) Ranking: A composite scoring system scores the $n+1$ code responses (including $y^{0}$ from origin dataset) and ranks them, identifying a winner $y^{w}$ and $n$ losers $\{y^{l_{i}}\}_{i=1}^{n}$. (3) Refining: Depending on the identity of the $y^{w}$, the selective critique strategy constructs a new single-round data $(x,y^{w})$ or two-round critique data $(x,y^{l_{1}},c,y^{w})$, where $y^{l_{1}}$ is the highest-scoring loser, $c$ is critique. The original data $(x,y)$ is replaced with new data and all the new data form a new dataset $D_{t+1}$. (4) Training: The new model $M_{t+1}$ is fine-tuned using $D_{t+1}$.
Distillation-based Code Fine-Tuning Unlike proprietary models, many studies focus on fine-tuning open-source code pre-trained models, which has greatly contributed to the rapid development of the code generation field. A key technique for achieving this is distilling data from teacher models. Code Alpaca Chaudhary (2023) introduces Self-Instruct Wang et al. (2023) to distill GPT-3. Following this, WizardCoder Luo et al. (2024) evolves more complex code instruction data using Evol-Instruct Xu et al. (2024). MagiCoder Wei et al. (2024b) proposes OSS-Instruct, where the model generates instructions and code responses sequentially based on open-source code snippets, thus creating more practical code data. In contrast to OSS-Instruct, InverseCoder Wu et al. (2024) reverses the order of instruction and code response generation. WaveCoder Yu et al. (2024) constructs a multi-task code dataset, enhancing the model’s versatility. Besides, OpenCodeInterpreter Zheng et al. (2024a) builds multi-turn code data, enabling the model to learn from execution feedback and human feedback. However, these methods aim to enable the model to learn by imitating the teacher, overlooking the potential for refinement through self-generated code.
Iterative Self-Refinement of LLMs Iterative self-refinement refers to LLMs enhancing themselves iteratively by utilizing self-generated responses with the help of external signals. One line of research Huang et al. (2023b); Madaan et al. (2023); Hu et al. (2024) focuses on enabling self-correction during the inference stage by iteratively calling LLMs and incorporating external signals. In the code domain, CodeT Chen et al. (2023), Self-Debugging Chen et al. (2024), and LDB Zhong et al. (2024) follow this approach. However, these prompt engineering methods with external feedback cannot improve the intrinsic capabilities. Another line of research focuses on iteratively training the model using self-generated outputs to enhance its intrinsic capabilities Dong et al. (2025); Yuan et al. (2024); Kim et al. (2025). These methods typically rely on preference learning, such as DPO Rafailov et al. (2023). CodeLutra Tao et al. (2024) has successfully applied this approach to the code domain, but it heavily depends on golden labels, which limits its applicability. In contrast to the works above, we propose the ACR method, which achieves iterative self-refinement by only using a simple Supervised Fine-Tuning (SFT) loss. This approach is orthogonal to prompt engineering methods like Self-Debugging and is more efficient and generalizable than CodeLutra.
3 Methodology
3.1 Overview
Adaptive Critique Refinement (ACR) aims to improve code LLMs by refining existing dataset with self-generated code and external critiques. The iterative application of ACR facilitates continuous improvement in code generation performance. As illustrated in Figure 2, ACR starts with an existing dataset and a code model fine-tuned on it. The method updates the existing dataset through a process of sampling, ranking, and refining. The updated dataset is used for the next round of fine-tuning, iteratively improving the model’s code generation capabilities.
In the following sections, we will introduce two key components of the aforementioned pipeline: the composite scoring system for ranking and the selective critiquing strategy for data refining. Finally, we will introduce the iterative training and the resulting RefineCoder model.
3.2 Composite Scoring System with LLM-as-a-Judge
A comprehensive and accurate evaluation of code response quality is the foundation for effective ACR. Previous works Chen et al. (2023); Zhang et al. (2024a) rely on generating test cases to evaluate code quality. However, this method has issues with inaccurate test case generation Jain et al. (2024), inability to handle special judge Quan et al. (2025), and failure to evaluate non-code content in response such as code comments and explanations.
Drawing on the judgment and evaluation abilities of LLMs Dong et al. (2025); Li et al. (2024), we employ an LLM-as-a-Judge as the backbone of the scoring system, supplemented by the Elo mechanism Elo and Sloan (1978) and a code executor to ensure effective evaluation. Specifically, given a programming question $x$ from the existing dataset $D_{t}$ and the corresponding code response list $\{y^{i}\}_{i=0}^{n}$, the Judge assigns pairwise relative scores for all code response pairs by identifying their differences and evaluating these discrepancies:
$\textrm{score}_{judge}^{ij},\ \textrm{score}_{judge}^{ji}=\textrm{Judge}(x,y^{i},y^{j}),$ (1)
where $i,j\in[0,n]$, $i\neq j$, and $\text{score}_{judge}^{ij}\in[1,10]$ denotes the relative score of the $i$-th response when compared to the $j$-th response. The prompt for LLM-as-a-Judge is shown in Figure 7. Next, we use Elo mechanism, a ranking algorithm that adjusts scores based on pairwise comparisons, to compute the pointwise score of each code response:
$\textrm{score}_{judge}^{i}=\textrm{Elo}\left(\left\{\textrm{score}_{judge}^{ij}\mid j\in[0,n],j\neq i\right\}\right),$ (2)
where $i\in[0,n]$ and $\textrm{score}_{judge}^{i}\in[0,1]$. For more details on Elo in the appendix A. In addition to the model scores provided by Judge and Elo, we also use a code executor to evaluate the executability of the code:
$\textrm{score}_{exec}^{i}=\textrm{Executor}(y^{i}),$ (3)
where $i\in[0,n]$ and $\text{score}_{exec}^{i}\in\{0,1\}$. Finally, we rank all code responses based on the composite score of them:
$\begin{split}&\textrm{score}^{i}=\textrm{score}^{i}_{judge}+\textrm{score}^{i}_{exec},\\ \end{split}$ (4)
$\begin{split}&y^{w},y^{l1},\dots,y^{l_{n}}=\textrm{Sort}(\{y^{i}\}_{i=0}^{n}\mid\{\textrm{score}^{i}\}_{i=0}^{n}),\end{split}$ (5)
where $\textrm{score}^{i}\in[0,2]$, $y^{w}$ and $y^{l_{1}}$ denote the code responses with the highest and second-highest scores, respectively.
3.3 Selective Critiquing Strategy with LLM-as-a-Critic
After obtaining the sorted code responses, we design the selective critiquing strategy as data construction engine of ACR. It effectively utilizes high-quality and low-quality self-generated code in distinct ways, while further providing external critiques for the low-quality data. This strategy resembles the process of a student first solving problems and then critiquing their answers based on provided solutions, rather than directly imitating the solutions Wang et al. (2025).
Concretely, when all self-generated code responses are of lower quality than the original code in the existing dataset, namely $y^{w}$ is $y^{0}$ and $y^{l_{1}}$ is the self-generated code, LLM-as-a-Critic is used to critique $y^{l_{1}}$, using $y^{0}$ as a reference (The prompt is shown in Figure 8):
$c=\textrm{Critic}(x,y^{w},y^{l_{1}}).$ (6)
Otherwise, if $y^{w}$ is the self-generated code response, we directly construct new single-round data. The data update rules are as follows:
$(x,y)\longrightarrow\begin{cases}(x,y^{l_{1}},c,y^{w}),&\text{if }y^{w}\text{ is }y^{0},\\ (x,y^{w}),&\text{otherwise}.\end{cases}$ (7)
If the data to be updated comes from a two-round critique rather than a single-round one, only the instruction and the final code response will be considered. Selective refining strategy updates all data to comprise a new instruction dataset $D_{t+1}$.
3.4 Iterative Training
The previous sections describes the process of a single ACR in Figure 2. Next, we describe the iterative ACR process and define RefineCoder:
$\circ$ RefineCoder Iter0 ($M_{0}$): The code model trained using the initial SFT dataset $D_{0}$.
$\circ$ RefineCoder Iter1 ($M_{1}$): The code model trained using dataset $D_{1}$, where $D_{1}$ is obtained by refining $D_{0}$ using $M_{0}$.
$\circ$ RefineCoder Itert ($M_{t}$): The code model trained using dataset $D_{t}$, where $D_{t}$ is obtained by refining $D_{t-1}$ using $M_{t-1}$.
We fine-tune the original base model from scratch in each iteration to prevent overfitting, consistent with previous iterative training works Zelikman et al. (2022); Wang et al. (2024a); Dong et al. (2025).
4 Experiments
4.1 Benchmarks
We use HumanEval (+) Chen et al. (2021) and MBPP(+) Austin et al. (2021); Liu et al. (2023) to assess the fundamental coding ability of code model. Meanwhile, we use LiveCodeBench Jain et al. (2024) and BigCodeBench-hard Zhuo et al. (2024) to measure the advanced coding ability of code model. We restrict data usage to the range from 240701 to 250201 of the LiveCodeBench to maintain its contamination-free nature.
Models Dataset HumanEval MBPP LCB BCB-hard Average
Base Plus Base Plus 2407-2502 Inst Comp
Proprietary Models
[…]
Claude-3.5-Sonnet-20240620 / 87.2 81.7 89.4 74.3 32.0 25.7 33.1 60.5
GPT-4o-20240806 / 92.7 87.2 87.6 72.2 30.5 25.0 36.5 61.7
Open-source Models 7B+ Scale
[…]
DeepSeek-V3 / 91.5 86.6 87.6 73.0 36.3 27.7 39.9 63.2
Llama3.3-70B-Instruct / 84.1 80.5 87.6 73.5 29.1 28.4 28.4 58.8
Open-source Models $\sim$7B Scale
[…]
DSCoder-6.7B-Instruct 2BT 74.4 71.3 74.9 65.6 12.8 10.1 15.5 46.4
MagiCoder-S-DS-6.7B 185K 76.8 71.3 79.4 69.0 13.4 13.5 12.8 48.0
OpenCodeInterpreter-DS-6.7B 178K 77.4 72.0 76.5 66.4 7.3 13.5 16.9 47.1
WaveCoder-Ultra-6.7B 130K 75.0 69.5 74.9 63.5 12.9 12.8 16.9 46.5
RefineCoder-DS-6.7B   Iter0 80K 73.8 67.7 77.5 65.1 13.6 18.2 10.1 46.6
RefineCoder-DS-6.7B   Iter1 80K 74.4 68.9 77.0 66.4 14.1 18.9 14.2 47.7+1.1
RefineCoder-DS-6.7B   Iter2 80K 74.4 70.3 79.6 66.9 14.3 19.6 14.2 48.5+1.9
RefineCoder-DS-6.7B   Iter3 80K 75.0 70.7 80.2 67.2 14.2 20.3 15.5 49.0+2.4
[Uncaptioned image]Qwen2.5-Coder-7B-Base / 61.6 53.0 76.9 62.9 / / 16.2      /
Qwen2.5-Coder-7B-Instruct Millions 88.4 84.1 83.5 71.7 18.1 20.3 20.3 55.2
RefineCoder-QW-7B   Iter0 20K 85.4 78.0 79.4 65.1 18.2 18.2 20.9 52.2
RefineCoder-QW-7B   Iter1 20K 86.0 79.3 84.7 71.4 17.4 16.2 18.9 53.4+1.2
RefineCoder-QW-7B   Iter2 20K 86.6 80.5 84.7 72.0 19.2 18.2 19.6 54.4+2.2
RefineCoder-QW-7B   Iter3 20K 87.2 81.1 85.2 73.0 19.1 19.6 20.9 55.2+3.0
Table 1: Main Results. LCB/BCB denotes LiveCodeBench/BigCodeBench. Inst/Comp denote Instruct/Complete. [Uncaptioned image] and [Uncaptioned image] denotes the two base models, and the rest are the instruction models. 2BT denotes 2B tokens.
4.2 Baselines
$\circ$ Proprietary Models: Gemini-1.5-Pro-002 Google (2024), Claude-3.5-Sonnet-20240620 Anthropic (2024), and GPT-4o-20240806 OpenAI (2024a).
$\circ$ Open-source Models 7B+ Scale: Qwen2.5-Coder-32B-Instruct Hui et al. (2024), DeepSeek-V3 DeepSeek-AI (2024), and Llama3.3-70B-Instruct Grattafiori et al. (2024).
$\circ$ Open-source Models $\sim$7B Scale: DSCoder-6.7B Series Guo et al. (2024), MagiCoder-S-DS-6.7B Wei et al. (2024b), OpenCodeInterpreter-DS-6.7B Zheng et al. (2024a), WaveCoder-Ultra-6.7B Yu et al. (2024), and Qwen2.5-Coder-7B Series Hui et al. (2024).
4.3 Initial SFT Dataset $D_{0}$
The existing datasets vary in quality and carry the risk of data leakage Wang et al. (2024b); Yu et al. (2024). To ensure the accuracy and reliability of the experimental results, we constructed a high-quality SFT dataset by combining the Evol-Instruct and SelfCodeAlign Wei et al. (2024a). Concretely, we first prompt GPT-4o to generate code-related concepts and corresponding programming instructions from Python code snippets. These code snippets have been preprocessed to ensure high quality and contamination-free. Then, we prompt GPT-4o again to iteratively evolve these instructions using Evol-Instruct strategy. Finally, we prompt GPT-4o to generate code responses for the instructions. Through this pipeline, we obtained an 80K high-quality and diverse python instruction dataset. See Appendix C for detailed data construction prompts.
Dataset Size HumanEval (+) MBPP (+) LCB BCB-hard
Magic-OSS-Instruct Wei et al. (2024b)  [^1] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K 75K 4.45 9.40 2.63 4.84
Magic-Evol-Instruct Wei et al. (2024b)  [^2] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K 110K 43.20 19.40 2.91 4.46
Evol-CodeAlpaca-v1 Luo et al. (2024)  [^3] https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1 110K 47.04 19.46 2.91 4.46
Code-FeedBack Zheng et al. (2024a)  [^4] https://huggingface.co/datasets/m-a-p/Code-Feedback 68K 30.50 17.67 3.16 4.78
Ours 80K 4.97 7.00 1.54 5.43
Table 2: The Test Leakage Indicator (TLI) quantifies data leakage by measuring the average maximum n-gram overlap between dataset samples and benchmark samples. The larger the TLI value, the more severe the data leakage. Values in red indicate severe data leakage.
4.4 Implement Details
Iterative Training We employ DSCoder-6.7B-Base and Qwen2.5-Coder-7B-Base as the base pre-trained models. To obtain an initial code model $M_{0}$, we fine-tune DSCoder-6.7B-Base on the 80K initial SFT dataset, resulting in RefineCoder-DS-6.7B (Iter0). Furthermore, since Qwen2.5-Coder-Base exhibits strong code understanding and generation capabilities, we accelerate our experimental iterations by randomly sampling 20K examples from the same dataset for fine-tune it, yielding RefineCoder-QW-7B (Iter0). The two RefineCoder models undergo three iterative training using their respective datasets under our ACR framework. In the ACR pipeline, the model self-samples 5 code responses with a temperature of 0.7. We use Qwen2.5-32B-Instruct Hui et al. (2024) as the Judge and Critic. We set the number of training epochs to 2, the global batch size to 64, and the learning rate to 5e-6, employing the AdamW optimizer with a cosine learning rate decay strategy. All our training is performed using 16 A100-80G GPUs, utilizing the LLaMA-Factory Zheng et al. (2024b).
Evaluation We use the Pass@1 metric to evaluate the performance of model on the benchmark. For baseline results, we prioritize those from the benchmark leaderboard  [^5] https://evalplus.github.io/leaderboard.html  [^6] https://livecodebench.github.io/leaderboard.html  [^7] https://huggingface.co/spaces/bigcode/bigcodebench-leaderboard or the original papers. If unavailable, we evaluate them locally using the same settings as the RefineCoder series.
4.5 Experimental Results
As shown in Table 1, the RefineCoder series achieves impressive performance gains on various code generation benchmarks. While the initial RefineCoder-DS-6.7B and RefineCoder-QW-7B exhibit moderate performance, the two models achieve an average Pass@1 improvement of 2.4 and 3.0, respectively, on all benchmarks after three iterations of refinement. Furthermore, it is worth noting that the RefineCoder series is fine-tuned with a very limited amount of data (80K or 20K). Despite this, the RefineCoder models (Iter3) still outperform or match baselines of the same size, which are fine-tuned with more extensive data.
Concretely, data refinement improves the fundamental programming abilities of the RefineCoder series, as demonstrated by continuous performance gains over iterations on HumanEval, HumanEval+, and MBPP+. Notably, RefineCoder-QW-7B shows a remarkable 7.9 point increase (Iter0 -> Iter3) in Pass@1 on MBPP+, highlighting the efficacy of this approach. In the more challenging benchmarks, LiveCodeBench and BigCodeBench-hard, the performance of the RefineCoder also improves after iterations. Compared to the baseline OpenCodeInterpreter, RefineCoder-DS-6.7B (Iter2) outperforms it by 7 points on LiveCodeBench, illustrating the strong competitive programming abilities of the RefineCoder series. These results demonstrate that the model can iteratively enhance code generation capabilities by self-generated code and critique with the help of our ACR method.
5 Further Study
5.1 Data Leakage Analysis
In Table 1, we observe that while the RefineCoder (Iter3) surpasses or matches the baselines on average pass@1, it still lags behind the state-of-the-art baseline of the same size on HumanEval (+) and MBPP+. To better explore this performance gap, we investigate potential data leakage in the datasets used by the baselines.
Concretely, we collect the open-source datasets used by MagiCoder-S-DS-6.7B, OpenCodeInterpreter-DS-6.7B, and WaveCoder-Ultra-6.7B (the datasets for the other baselines are not public) and evaluate data leakage using the TLI metric proposed by Wang et al. (2024b). The results are shown in Table 2, Magic-OSS-Instruct and our dataset are entirely free from data leakage. In contrast, Magic-Evol-Instruct, Evol-CodeAlpaca-v1, and Code-FeedBack demonstrate severe data leakage with respect to HumanEval (+) and MBPP (+). These datasets are used by the three aforementioned baselines, making it challenging for RefineCoder to surpass these baselines. We also present a intuitive scatter plot of similarity scores in the appendix E for a more detailed analysis.
HumanEval+ (%) Best Worst Error Rate
RefineCoder-DS-6.7B Iter3 76.2 60.4-15.8 4.3
RefineCoder-QW-7B Iter3 81.7 75.0-6.7 3.0
MBPP+ (%) Best Worst Error Rate
RefineCoder-DS-6.7B Iter3 65.3 60.1-5.2 6.3
RefineCoder-QW-7B Iter3 70.4 69.6-0.8 5.8
Table 3: Best and Worst denote the performance of code with the highest and lowest scores. Error rate denotes the proportion of cases where the best code fails the test cases, but the worst code passes.
5.2 The effectiveness of Composite Scoring System
The performance improvement through iterative training preliminarily validates the effectiveness of the scoring system. In this section, we further verify it by enabling the model sampling 10 code responses for each programming question in benchmarks and scoring them. As shown in Table 3, the performance of the highest-scoring code far exceeds that of the lowest-scoring code, while the scoring system maintains a low error rate. This clearly validates its effectiveness.
5.3 The effectiveness of Selective Critique Strategy
To validate the effectiveness of the selective critique strategy in ACR, we conduct ablation experiments. As shown in Figure 3, for each iteration, we construct two variants of the RefineCoder without the 3.1 module in Figure 2, and another without the 3.2 module. During the iteration process, ablating either module leads to a performance drop in both RefineCoder-QW-7B and RefineCoder-DS-6.7B, indicating the effectiveness of our strategy. Furthermore, removing the second-round critique data has a greater impact on performance. We believe this action prevents the model from reflecting on self-generated low-quality codes and external critiques, thereby hindering performance improvement.
Figure 3: Ablation Study. The y-axis denotes the average pass@1 value on all benchmarks.
5.4 Multilingual Code Generation
Figure 4: Evaluation with feedback using RefineCoder-QW-7B.
Models C++ C# Java Bash TypeScript JavaScript AVG
RefineCoder-DS-6.7B   Iter0 56.5 58.9 58.9 40.5 62.3 64.6 57.2
RefineCoder-DS-6.7B   Iter1 62.1 61.4 55.1 35.4 65.4 66.5 58.2
RefineCoder-DS-6.7B   Iter2 63.4 56.2 56.2 38.6 65.4 65.8 57.5
RefineCoder-DS-6.7B   Iter3 63.4 58.2 57.0 39.2 62.9 65.8 57.8
RefineCoder-QW-7B   Iter0 60.2 73.4 68.4 49.4 78.0 78.3 68.7
RefineCoder-QW-7B   Iter1 62.7 74.1 65.2 51.3 76.1 77.0 68.6
RefineCoder-QW-7B   Iter2 63.8 74.1 69.6 48.7 78.0 77.6 69.4
RefineCoder-QW-7B   Iter3 60.2 75.3 73.4 53.2 79.2 77.6 70.6
Table 4: Performance of RefineCoder on the MultiPL-E.
As shown in Table 4, we also evaluate the out-of-distribution (OOD) code generation performance on multilingual benchmark MultiPL-E Cassano et al. (2022), despite fine-tuning on a Python-only dataset. After one and three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve optimal average performance, surpassing the initial model by 1.0 and 1.9 points, respectively. In particular, RefineCoder-QW-7B (Iter3) achieves best results among half of the programming languages. This demonstrates that the effectiveness of our ACR method generalizes well to the OOD code generation domain.
5.5 Evaluation with External Feedback
Iterative ACR not only improves the one-off code generation performance of RefineCoder but also endows it with the ability to correct errors based on feedback. Following  Zheng et al. (2024a), we design two types of external feedback to evaluate this ability of RefineCoder: 1) Execution Feedback: Model refines its self-generated erroneous code based on execution results from the executor. 2) Human Feedback: GPT-4o first analyzes the programming question, initial error code, and execution results to generate improvement suggestions that mimic human thinking. Model then refines the code based on these suggestions. The results of RefineCoder-QW-7B on HumanEval (+) and MBPP (+) are shown in Figure 4. It can be observed that external feedback improves performance on all benchmarks, with human feedback yielding a more significant enhancement. Surprisingly, with the help of human-like improvement suggestions, RefineCoder-QW-7B (Iter1 and Iter3) outperforms GPT-4o on three benchmarks, demonstrating its ability to understand feedback and refine code.
Models HumanEval MBPP
Base Plus Base Plus
Qwen2.5-Coder-7B-Base 61.6 53.0 76.9 62.9
     +DPO 59.8 52.4 44.2 32.6
     +ORPO 85.4 79.9 84.4 70.4
RefineCoder-QW-7B Iter0 85.4 78.0 79.4 65.1
     +DPO 61.0 50.0 60.6 42.3
     +ORPO 86.6 79.3 81.7 69.3
Table 5: Performance comparison of different preference optimization strategies in reinforcement learning.
5.6 Preference Optimization
During the iterative process, a substantial amount of preference data is generated. For each programming instruction, we select the top-1 and bottom-1 ranked responses to form preference pairs, thereby exploring the potential of preference learning. As shown in Table 5, we use DPO Rafailov et al. (2023) and ORPO Hong et al. (2024) to optimize the base and SFT models. We observe the opposite results due to the differences in preference strategies. DPO leads to a significant performance decline due to the intrinsic risk of reducing the likelihood of correct code during training Feng et al. (2024); Tao et al. (2024); Pal et al. (2024). In contrast, ORPO builds upon the SFT loss to maximize the likelihood of correct code, resulting in improved performance. The results are consistent with those presented in the CodeLutra Tao et al. (2024), reflecting the indispensability of SFT loss and the critical impact of preference strategy design on performance.
6 Conclusions
In this paper, we propose the Adaptive Critique Refinement (ACR) method to iteratively refine code LLMs using self-generated code and external critique, which breaks away from the traditional teacher distillation paradigm and improves the intrinsic one-off code generation ability. We design a composite scoring system and a selective critiquing strategy. These two components are centred around LLM-as-a-Judge and LLM-as-a-Critic to evaluate and critique self-generated code. This simulates the process where a student solves a problem independently and then refines it by comparing it with the reference answer. We develop the RefineCoder-DS-6.7B and RefineCoder-QW-7B models and demonstrate the effectiveness of iterative ACR on HumanEval (+), MBPP (+), LiveCodeBench, and BigCodeBench-hard. Further studies reveal the impact of the components in ACR and demonstrate its OOD code generation ability.
7 Limitations
Although Adaptive Critique Refinement (ACR) and the associated RefineCoder demonstrate significant improvements in the code generation task, several limitations remain. First, ACR is primarily designed to refine code responses based on programming instructions, necessitating an initial set of high-quality and diverse instructions. Therefore, a specialized code instruction generator is still required to make ACR more automated. Furthermore, while ACR can apply to other reasoning-intensive tasks, such as mathematics, this paper has not fully explored these domains.
Appendix A Elo Algorithm
Firstly, we initialize Elo score $E^{i}=1000$ for each code response $y^{i}$. Then, we iteratively update the Elo scores by using the relative scores between any two responses. Taking the response pair $y^{i}$ and $y^{j}$ as an example, we obtain their match results $S^{i}$ and $S^{j}$:
$S^{i}=\frac{\textrm{score}_{judge}^{ij}}{\textrm{score}_{judge}^{ij}+\textrm{score}_{judge}^{ji}},$ (8)
$S^{j}=\frac{\textrm{score}_{judge}^{ji}}{\textrm{score}_{judge}^{ij}+\textrm{score}_{judge}^{ji}}.$ (9)
Then we obtain the expected probabilities of winning $P^{ij}$ and $P^{ji}$ for $y^{i}$ and $y^{j}$:
$P^{ij}=\frac{1}{1+10^{\frac{E^{j}-E^{i}}{400}}},$ (10)
$P^{ji}=1-P^{ij}.$ (11)
Finally, we update the Elo score:
$E^{i}_{new}=E^{i}+K\times(S^{i}-P^{ij}),$ (12)
$E^{j}_{new}=E^{j}+K\times(S^{j}-P^{ji}).$ (13)
where $K=32$ is a adjustment factor. After the iteration, we obtain the Elo score for each code response and then normalize it to derive the final judge score:
$\textrm{score}^{i}_{judge}=\frac{E_{i}-\min(E)}{\max(E)-\min(E)},$ (14)
where $i\in[0,n]$ and $\textrm{score}^{i}_{judge}\in[0,1]$.
Appendix B Prompts for Judge and Critic
The prompt for pairwise LLM-as-a-Judge and LLM-as-a-Critic are shown in Figure 7 and  8.
Appendix C Prompts for constructing SFT Dataset
We called GPT-4o three times to create the SFT dataset, with the following prompts used:
$\circ$ Prompt for generating code-related concepts and programming instructions, as shown in Figure 9.
$\circ$ Prompts for evolving existing instructions, as shown in Figures 10,  11,  1213141516,  17,  18.
$\circ$ For generating code responses, We directly feed the instructions to the model.
Appendix D Prompts for Evaluation
For Humaneval (+) and MBPP (+), we use the prompts designed by OpenCodeInterpreter Zheng et al. (2024a); for Livecodebench and Bigcodebench-hard, we use the official prompts. When conducting evaluations with execution feedback, we have designed a prompt as illustrated in Figure 19. When executing human feedback, following OpenCodeInterpreter, we generate two prompts: one for generating improvement suggestions and the other for refine code according to human feedback, as shown in Figures 21 and  20.
Appendix E Data Leakage Analysis
We use 4-grams and 5-grams to compute the similarity between the instructions in the training set and the questions in the benchmark. And then we draw scatter plots by selecting the top-200 data points with the highest similarity from each dataset, as shown in Figure 5. The scatter plot visually illustrates the data leakage phenomenon in the dataset, primarily concentrated in the HumanEval (+) and the latter half of the MBPP (+). This makes it difficult for our model to surpass baselines on these two benchmarks. In contrast, on the leakage-free LiveCodeBench and BigCodeBench-hard benchmarks, our model can outperform all baselines.
Appendix F Scaling Law of Iterative ACR
We study the performance trends of RefineCoder-DS-6.7B obtained using ACR and the standard SFT method as the dataset size increases. For the SFT method, we directly trained the model using four datasets of different sizes. For RefineCoder series, we employ an iterative training strategy. For example, after obtaining RefineCoder-DS-10k by fing-tuning on the 10k dataset, we apply ACR to update this 10k dataset, and then randomly sample another 10k from the data pool to form a 20k dataset. The model is fine-tuned again to obtain RefineCoder-DS-20k. The experimental results are shown in Figure 6, as the dataset size increases, our model achieves greater performance improvements compared to the SFT model on HumanEval and MBPP. Moreover, the performance gap between the two models does not show any signs of narrowing, underscoring the impressive scaling capabilities of the iterative ACR method.
Figure 5: Scatter plot of similarity scores between the datasets and four benchmarks, with average similarity scores for different datasets indicated by horizontal dashed lines. We selected the top 200 data points with the highest similarity from each dataset. Different y-axis ranges are set for better visualization.
Figure 6: Performance curves of RefineCoder obtained using iterative ACR and the standard SFT model as the dataset size increases.
Figure 7: The prompt template for pairwise LLM-as-a-Judge.
Figure 8: The prompt template for pairwise LLM-as-a-Critic.
Figure 9: The prompt template for generating code-related concepts and instructions.
Figure 10: The prompt template for Addition evolution.
Figure 11: The prompt template for Breath evolution.
Figure 12: The prompt template for Complexity evolution.
Figure 13: The prompt template for Concretizing evolution.
Figure 14: The prompt template for Deepening evolution.
Figure 15: The prompt template for Diversion evolution.
Figure 16: The prompt template for Increase evolution.
Figure 17: The prompt template for Misdirection evolution.
Figure 18: The prompt template for Reason evolution.
Figure 19: The prompt template for Execution Feedback.
Figure 20: The prompt template for Human Feedback.
Figure 21: The prompt template for generating improvement suggestions.

### v2

@@
-RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation
+RefineCoder: Iterative Improving Large Language Models through Adaptive Critique Refinement for Code Generation
 Abstract
 Code generation has attracted increasing attention with the rise of Large Language Models (LLMs).
-Many studies have developed powerful code LLMs by synthesizing code-related instruction data and applying supervised fine-tuning.
+Many studies have developed powerful code LLMs by synthesizing code-related instruction data and applying Supervised Fine-Tuning (SFT).
 However, these methods are limited by teacher model distillation and ignore the potential of iterative refinement by self-generated code.
-In this paper, we propose Adaptive Critique Refinement (ACR), which enables the model to refine itself by self-generated code and external critique, rather than directly imitating the code responses of the teacher model.
-Concretely, ACR includes a composite scoring system with LLM-as-a-Judge to evaluate the quality of code responses and a selective critique strategy with LLM-as-a-Critic to critique self-generated low-quality code responses.
+In this paper, we propose Adaptive Critique Refinement (ACR), which enables the model to refine itself by self-generated code response and teacher-generated critique, rather than directly imitating the teacher-generated code responses.
+Concretely, ACR includes a composite scoring system with LLM-as-a-Judge to evaluate the quality of code responses and a selective critiquing strategy with LLM-as-a-Critic to critique self-generated low-quality code responses.
 We develop the RefineCoder series by iteratively applying ACR, achieving continuous performance improvement on multiple code generation benchmarks.
-Compared to the baselines of the same size, our proposed RefineCoder series can achieve comparable or even superior performance using less data.
-RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation
+Compared to the baselines of the same size, our proposed RefineCoder series can achieve superior performance using less data.
+RefineCoder: Iterative Improving Large Language Models through Adaptive Critique Refinement for Code Generation
 Changzhi Zhou1thanks:   Work done during internship at Meituan.
@@
 (2024).
-These LLMs demonstrate strong capabilities in code understanding owing to their pre-training on large-scale code corpus.
-However, they must undergo post-training to become user-friendly and effective for code-related tasks.
-Figure 1: Comparison of two fine-tuning paradigms. DFT uses code distilled by teacher model, whereas ACR adaptively constructs distinct data formats infusing self-generated code and external critique.
+Despite their strong capabilities in code understanding, these LLMs typically require additional post-training to become effective and user-friendly for code-related tasks.
+Figure 1: Comparison of two fine-tuning paradigms.
+DFT uses code distilled by teacher model, whereas ACR adaptively constructs distinct data formats infusing self-generated code and teacher-generated critique.
 As a popular post-training technology, Distillation-based Fine-Tuning (DFT) leverages code instruction data synthesized by powerful teacher models (often proprietary LLMs like GPT-4) to fine-tune code LLMs, as shown in Figure 1.
@@
 (2024b) based on this paradigm have been proposed recently.
-For example, Code Alpaca Chaudhary (2023) initially establishes this synthesis paradigm using Self-Instruct Wang et al.
+For example, Code Alpaca Chaudhary (2023) first establishes this synthesis paradigm using Self-Instruct Wang et al.
 (2023).
@@
 (2024) prompts LLMs to evolve more intricate code instructions, while OSS-Instruct Wei et al.
-(2024b) generates more realistic code instructions based on open-source code snippets from The Stack Kocetkov et al.
+(2024b) generates more realistic code instructions based on open-source code snippets Kocetkov et al.
 (2022).
@@
 However, this paradigm of teacher model distillation inevitably suffers from a critical limitation: the performance of the student model largely relies on the teacher model, ignoring the potential of iterative refinement by self-generated code.
-Apart from the research on code instruction fine-tuning, some studies Chen et al.
-(2023); Huang et al.
-(2023a); Chen et al.
-(2024, 2025) have explored the ability of LLMs to refine self-generated code during inference.
+Apart from research on code instruction fine-tuning, some other studies Chen et al.
+(2023, 2024, 2025); Jiang et al.
+(2025) have explored the ability of LLMs to refine self-generated code during inference.
 For example, Self-Debugging Chen et al.
@@
 They cannot improve the intrinsic one-off code generation ability of LLMs.
-Besides, multiple calls to LLMs increase the inference latency.
-In this paper, we propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm to enhance the intrinsic one-shot code generation ability of LLMs, and on the basis of which we develop a series of code LLMs named RefineCoder.
-ACR goes beyond traditional teacher distillation by using both self-generated code and external critique to refine the model.
-Specifically, ACR tailors the distinct new fine-tuning dataset by first scoring and critiquing the self-generated multiple code responses, and then adaptively forms new samples according to the scoring and critiquing results.
-The two types of samples are shown in Figure 1.
-Based on the new dataset, we fine-tune RefineCoder models.
-Finally, we iteratively applying the above ACR process to achieve continuous improvement in code generation capabilities of RefineCoder.
-Notably, in ACR, we design a composite scoring system and a selective critiquing strategy to better score code responses and selectively critique low-quality code.
-The former combines LLM-as-a-Judge, an Elo rating mechanism Elo and Sloan (1978), and a code executor for accurate and comprehensive evaluation.
-The latter adaptively constructs two types of data by comparing the scores of self-generated code with those of original code.
-When self-generated code performs worse than the original code, an LLM-as-a-Critic is introduced to provide specific critiques.
+Besides, multiple calls to LLMs increase inference latency.
+In this paper, we propose a novel fine-tuning paradigm called Adaptive Critique Refinement (ACR).
+It evolves through a combination of self-generated code and teacher-driven critique, thereby enhancing its capacity for one-shot code generation and surpassing the limitations of conventional teacher distillation methods.
+Specifically, as shown in Figure 1, ACR evaluates both self-generated and teacher-generated code, adaptively critiques low-quality self-generated responses, and generates new samples based on the evaluation and critique results.
+This process is analogous to a student initially attempting to solve problems independently and later comparing their solutions against given answers, rather than simply copying the provided solutions.
+By iteratively applying the ACR process, the model achieves continuous improvement in its code generation capabilities.
+Notably, we design a composite scoring system and a selective critique strategy to perform scoring and critique, respectively.
+The scoring system combines LLM-as-a-Judge and a code executor to provide accurate and comprehensive evaluations.
+The critique strategy adaptively constructs two types of data and incorporates LLM-as-a-Critic.
 We conduct iterative ACR based on DS-Coder-6.7B-Base Guo et al.
 (2024) and Qwen2.5-Coder-7B-Base Hui et al.
-(2024), resulting in the RefineCoder series.
-After three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve average pass@1 improvements of 2.4 and 3.0, respectively, on the HumanEval (+), MBPP (+), LiveCodeBench, and the BigCodeBench-hard benchmark Chen et al.
+(2024), leading to the development of the RefineCoder series.
+After three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B demonstrate significant improvements in average pass@1 performance on the HumanEval(+), MBPP(+), LiveCodeBench, and BigCodeBench (hard) Chen et al.
 (2021); Austin et al.
@@
 The key contributions of this paper are as follows:
-1) We propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm that refines code LLMs with self-generated code and external critique, on the basis of which we develop a series of strong code LLMs (RefineCoder).
-2) To ensure the efficacy of ACR, we design a composite scoring system with LLM-as-a-Judge and a selective critique strategy with LLM-as-a-Critic to score and critique self-generated code.
-3) Experimental results from the RefineCoder series show that iterative ACR continuously enhances code generation performance.
-After three iterations, the RefineCoder series achieves comparable or even superior performance than baselines of the same size while requiring less data.
+1) We propose Adaptive Critique Refinement (ACR), a novel fine-tuning paradigm that refines code LLMs using self-generated code and external critique, on the basis of which we develop a series of strong code LLMs namely RefineCoder.
+2) To ensure the effectiveness of ACR, we design a composite scoring system with LLM-as-a-Judge and a selective critiquing strategy with LLM-as-a-Critic, enabling both accurate scoring and critique of code responses.
+3) Experimental results from the RefineCoder series show that iterative ACR continuously improves code generation performance.
+After three iterations, the RefineCoder models outperform baselines of the same size while requiring less data.
 2 Related Work
@@
 (1) Sampling: The model $M_{t}$ samples $n$ code responses $\{y^{i}\}_{i=1}^{n}$ with high temperature for a programming instruction $x$ in the dataset $D_{t}$.
-(2) Ranking: A composite scoring system scores the $n+1$ code responses (including $y^{0}$ from origin dataset) and ranks them, identifying a winner $y^{w}$ and $n$ losers $\{y^{l_{i}}\}_{i=1}^{n}$.
-(3) Refining: Depending on the identity of the $y^{w}$, the selective critique strategy constructs a new single-round data $(x,y^{w})$ or two-round critique data $(x,y^{l_{1}},c,y^{w})$, where $y^{l_{1}}$ is the highest-scoring loser, $c$ is critique.
+(2) Ranking: A composite scoring system first scores all self-generated code responses $\{y^{i}\}_{i=1}^{n}$, then selects the highest-scoring response for pairwise comparison with the original code response $y^{0}$ to identify the winner $y^{w}$ and the loser $y^{l_{1}}$.
+(3) Refining: Depending on the identity of the $y^{w}$, the selective critiquing strategy constructs a new single-turn data $(x,y^{w})$ or two-turn critique data $(x,y^{l_{1}},c,y^{w})$, where $c$ is critique.
 The original data $(x,y)$ is replaced with new data and all the new data form a new dataset $D_{t+1}$.
@@
 One line of research Huang et al.
-(2023b); Madaan et al.
+(2023); Madaan et al.
 (2023); Hu et al.
@@
 (2025); Yuan et al.
+(2024); Yan et al.
 (2025); Kim et al.
@@
 In contrast to the works above, we propose the ACR method, which achieves iterative self-refinement by only using a simple Supervised Fine-Tuning (SFT) loss.
-This approach is orthogonal to prompt engineering methods like Self-Debugging and is more efficient and generalizable than CodeLutra.
+Our approach is orthogonal to prompt engineering methods like Self-Debugging and is more efficient and generalizable than CodeLutra.
 3 Methodology
 3.1 Overview
-Adaptive Critique Refinement (ACR) aims to improve code LLMs by refining existing dataset with self-generated code and external critiques.
+Adaptive Critique Refinement (ACR) aims to improve code LLMs by refining the existing dataset with self-generated code and external critique.
 The iterative application of ACR facilitates continuous improvement in code generation performance.
-As illustrated in Figure 2, ACR starts with an existing dataset and a code model fine-tuned on it.
-The method updates the existing dataset through a process of sampling, ranking, and refining.
+As illustrated in Figure 2, ACR starts with a current dataset and a code model fine-tuned on it.
+The method updates this dataset through a process of sampling, ranking, and refining.
 The updated dataset is used for the next round of fine-tuning, iteratively improving the model’s code generation capabilities.
-In the following sections, we will introduce two key components of the aforementioned pipeline: the composite scoring system for ranking and the selective critiquing strategy for data refining.
-Finally, we will introduce the iterative training and the resulting RefineCoder model.
+In the following sections, we will introduce two key components of the pipeline above: the composite scoring system for ranking and the selective critiquing strategy for data refining.
+Finally we describe the iterative training and the resulting RefineCoder model.
 3.2 Composite Scoring System with LLM-as-a-Judge
@@
 (2025); Li et al.
-(2024), we employ an LLM-as-a-Judge as the backbone of the scoring system, supplemented by the Elo mechanism Elo and Sloan (1978) and a code executor to ensure effective evaluation.
-Specifically, given a programming question $x$ from the existing dataset $D_{t}$ and the corresponding code response list $\{y^{i}\}_{i=0}^{n}$, the Judge assigns pairwise relative scores for all code response pairs by identifying their differences and evaluating these discrepancies:
-$\textrm{score}_{judge}^{ij},\ \textrm{score}_{judge}^{ji}=\textrm{Judge}(x,y^{i},y^{j}),$ (1)
-where $i,j\in[0,n]$, $i\neq j$, and $\text{score}_{judge}^{ij}\in[1,10]$ denotes the relative score of the $i$-th response when compared to the $j$-th response.
-The prompt for LLM-as-a-Judge is shown in Figure 7.
-Next, we use Elo mechanism, a ranking algorithm that adjusts scores based on pairwise comparisons, to compute the pointwise score of each code response:
-$\textrm{score}_{judge}^{i}=\textrm{Elo}\left(\left\{\textrm{score}_{judge}^{ij}\mid j\in[0,n],j\neq i\right\}\right),$ (2)
-where $i\in[0,n]$ and $\textrm{score}_{judge}^{i}\in[0,1]$.
-For more details on Elo in the appendix A.
-In addition to the model scores provided by Judge and Elo, we also use a code executor to evaluate the executability of the code:
-$\textrm{score}_{exec}^{i}=\textrm{Executor}(y^{i}),$ (3)
-where $i\in[0,n]$ and $\text{score}_{exec}^{i}\in\{0,1\}$. Finally, we rank all code responses based on the composite score of them:
-$\begin{split}&\textrm{score}^{i}=\textrm{score}^{i}_{judge}+\textrm{score}^{i}_{exec},\\ \end{split}$ (4)
-$\begin{split}&y^{w},y^{l1},\dots,y^{l_{n}}=\textrm{Sort}(\{y^{i}\}_{i=0}^{n}\mid\{\textrm{score}^{i}\}_{i=0}^{n}),\end{split}$ (5)
-where $\textrm{score}^{i}\in[0,2]$, $y^{w}$ and $y^{l_{1}}$ denote the code responses with the highest and second-highest scores, respectively.
+(2024), we employ LLM-as-a-Judge as the backbone of the scoring system, supplemented by a code executor to ensure effective evaluation.
+Specifically, we perform a multi-faceted, point-wise judgment of each self-generated code response $y^{i}$ ($1\leq i\leq n$):
+$s^{i}_{1},s^{i}_{2},...,s^{i}_{m}=\textrm{Multi-Faceted Judge}(x,y^{i}),$ (1)
+where $s^{i}_{j}$ denotes the score for $y^{i}$ under the $j$-th evaluation criterion, and $m$ is the total number of criteria (e.g., code correctness, clarity).
+Next, we calculate the final point-wise score for each self-generated code response by computing the weighted sum and normalizing the result:
+$s^{i}$ $=\sum_{j=1}^{m}\alpha_{j}s_{j}^{i},$ (2)
+$\text{score}^{i}_{\text{point}}$ $=\frac{s^{i}-s_{\min}}{s_{\max}-s_{\min}},$ (3)
+where $\alpha_{j}$ is the weight of the $j$-th criterion, and $s_{\min}$ and $s_{\max}$ are the minimum and maximum scores across self-generated code responses.
+In addition, we use a code executor to evaluate the executability of all code responses (including $y^{0}$):
+$\textrm{score}_{\text{exec}}^{i}=\textrm{Executor}(y^{i}),$ (4)
+where $i\in[0,n]$ and $\textrm{score}_{\text{exec}}^{i}\in\{0,1\}$. The highest-scoring self-generated code is obtained by summing the point-wise score and execution score:
+$y^{self}=\arg\max_{i}(\text{score}^{i}_{\text{point}}+\text{score}^{i}_{\text{exec}}).$ (5)
+where $i\in[1,n]$. Finally, we determine the winner between the best self-generated code $y^{self}$ and the teacher-generated code $y^{0}$ based on the pairwise judgment and execution scores:
+$\textrm{score}_{\text{pair}}^{self},\textrm{score}_{\text{pair}}^{0}=\textrm{PairJudge}(x,y^{self},y^{0}),$ (6)
+$\textrm{score}^{i}=\textrm{score}^{i}_{\text{pair}}+\textrm{score}^{i}_{\text{exec}},$ (7)
+$y^{w},y^{l_{1}}=\begin{cases}y^{self},y^{0},&\text{if }\textrm{score}^{self}>\textrm{score}^{0},\\ y^{0},y^{self},&\text{if }\textrm{score}^{0}>\textrm{score}^{self},\end{cases}$ (8)
+where $i\in\{self,0\}$, $\textrm{score}^{i}_{pair}\in[0,1]$, and $y^{w}$ and $y^{l_{1}}$ denote the winner and loser, respectively.
+The prompts for point-wise and pair-wise judgment are shown in Figures 8 and 9.
 3.3 Selective Critiquing Strategy with LLM-as-a-Critic
-After obtaining the sorted code responses, we design the selective critiquing strategy as data construction engine of ACR.
-It effectively utilizes high-quality and low-quality self-generated code in distinct ways, while further providing external critiques for the low-quality data.
-This strategy resembles the process of a student first solving problems and then critiquing their answers based on provided solutions, rather than directly imitating the solutions Wang et al.
+After selecting winner and loser based on quality, we design the selective critiquing strategy as the data construction engine of ACR.
+It effectively utilizes high-quality and low-quality self-generated code in distinct ways, while further providing external critique for the low-quality data.
+This strategy mirrors the process of a student first solving problems and then critiquing their answers based on provided solutions rather than directly imitating the solutions Wang et al.
 (2025).
-Concretely, when all self-generated code responses are of lower quality than the original code in the existing dataset, namely $y^{w}$ is $y^{0}$ and $y^{l_{1}}$ is the self-generated code, LLM-as-a-Critic is used to critique $y^{l_{1}}$, using $y^{0}$ as a reference (The prompt is shown in Figure 8):
-$c=\textrm{Critic}(x,y^{w},y^{l_{1}}).$ (6)
-Otherwise, if $y^{w}$ is the self-generated code response, we directly construct new single-round data. The data update rules are as follows:
+Concretely, when all self-generated code responses are of lower quality than the original code in the current dataset (i.e., $y^{w}$ is $y^{0}$ from origin dataset and $y^{l_{1}}$ is the best self-generated code response), LLM-as-a-Critic is used to critique $y^{l_{1}}$, using $y^{0}$ as a reference (The prompt is shown in Figure 10):
+$c=\textrm{Critic}(x,y^{w},y^{l_{1}}),$ (9)
+where $c$ is the critique of $y^{l_{1}}$. Otherwise, if $y^{w}$ is the self-generated code response, we directly construct new single-turn data. The data update rules are as follows:
 $(x,y)\longrightarrow\begin{cases}(x,y^{l_{1}},c,y^{w}),&\text{if }y^{w}\text{ is }y^{0},\\ (x,y^{w}),&\text{otherwise}.\end{cases}$ (10)
-If the data to be updated comes from a two-round critique rather than a single-round one, only the instruction and the final code response will be considered.
-Selective refining strategy updates all data to comprise a new instruction dataset $D_{t+1}$.
+If the data to be updated comes from a two-turn data rather than a single-turn one, only the instruction and the final code response will be considered.
+The selective critiquing strategy updates all data to comprise a new instruction dataset $D_{t+1}$.
 3.4 Iterative Training
-The previous sections describes the process of a single ACR in Figure 2. Next, we describe the iterative ACR process and define RefineCoder:
-$\circ$ RefineCoder Iter0 ($M_{0}$): The code model trained using the initial SFT dataset $D_{0}$.
-$\circ$ RefineCoder Iter1 ($M_{1}$): The code model trained using dataset $D_{1}$, where $D_{1}$ is obtained by refining $D_{0}$ using $M_{0}$.
-$\circ$ RefineCoder Itert ($M_{t}$): The code model trained using dataset $D_{t}$, where $D_{t}$ is obtained by refining $D_{t-1}$ using $M_{t-1}$.
-We fine-tune the original base model from scratch in each iteration to prevent overfitting, consistent with previous iterative training works Zelikman et al.
-(2022); Wang et al.
-(2024a); Dong et al.
-(2025).
+The previous sections describe the process of a single ACR in Figure 2. Next, we define the RefineCoder in the iterative ACR process:
+$\circ$ RefineCoder Iter0 ($M_{0}$): The code model is trained using the initial SFT dataset $D_{0}$.
+$\circ$ RefineCoder Iter1 ($M_{1}$): The code model is trained using dataset $D_{1}$, where $D_{1}$ is obtained by refining $D_{0}$ using $M_{0}$.
+$\circ$ RefineCoder Itert ($M_{t}$): The code model is trained using dataset $D_{t}$, where $D_{t}$ is obtained by refining $D_{t-1}$ using $M_{t-1}$.
+We fine-tune the original pre-trained model from scratch in each iteration to prevent overfitting, consistent with previous iterative training works Wang et al. (2024a); Dong et al. (2025).
+Models Data Size LiveCodeBench BigCodeBench-hard Average
+v3-v5 v5-v6 Instruct Complete
+[Uncaptioned image] DSCoder-6.7B-Base / / / / 13.5 /
+DSCoder-6.7B-Instruct 2BT 12.8 14.0 10.1 15.5 13.1
+MagiCoder-S-DS-6.7B 185K 13.4 14.6 13.5 12.8 13.6
+OpenCodeInterpreter-DS-6.7B 178K 7.3 15.2 13.5 16.9 13.2
+WaveCoder-Ultra-6.7B 130K 12.9 14.5 12.8 16.9 14.3
+RefineCoder-DS-6.7B   Iter0 80K 13.1 15.5 18.2 10.1 14.2
+RefineCoder-DS-6.7B   Iter1 80K 14.3 17.2 18.9 14.2 16.2↑2.0
+RefineCoder-DS-6.7B   Iter2 80K 14.3 17.8 19.6 14.2 16.5↑2.3
+RefineCoder-DS-6.7B   Iter3 80K 14.2 17.9 20.1 15.5 16.9↑2.7
+[Uncaptioned image]Qwen2.5-Coder-7B-Base / / / / 16.2    /
+Qwen2.5-Coder-7B-Instruct Millions 18.1 18.9 20.3 20.3 19.4
+RefineCoder-QW-7B   Iter0 80K 18.0 18.7 18.5 19.3 18.6
+RefineCoder-QW-7B   Iter1 80K 19.6 21.3 20.7 21.8 20.9↑2.3
+RefineCoder-QW-7B   Iter2 80K 20.5 21.5 20.6 22.9 21.4↑2.8
+RefineCoder-QW-7B   Iter3 80K 21.7 21.9 20.8 22.3 21.7↑3.1
+Table 1: Results on LiveCodeBench and BigCodeBench-hard.
+v3-v5 and v5-v6 refer to the questions from the periods 202407-202502 and 202502-202505, respectively.
+[Uncaptioned image] and [Uncaptioned image] denotes the two base models, and the rest are the instruction models.
+2BT denotes 2B tokens.
 4 Experiments
@@
 We use HumanEval(+) Chen et al.
-(2021) and MBPP(+) Austin et al.
+(2021), MBPP(+) Austin et al.
 (2021); Liu et al.
+(2023) and BigCodeBench Zhuo et al.
 (2024) to assess the fundamental coding ability of code model.
 Meanwhile, we use LiveCodeBench Jain et al.
-(2024) and BigCodeBench-hard Zhuo et al.
-(2024) to measure the advanced coding ability of code model.
-We restrict data usage to the range from 240701 to 250201 of the LiveCodeBench to maintain its contamination-free nature.
-Models Dataset HumanEval MBPP LCB BCB-hard Average
-Base Plus Base Plus 2407-2502 Inst Comp
-Proprietary Models
-[…]
-Claude-3.5-Sonnet-20240620 / 87.2 81.7 89.4 74.3 32.0 25.7 33.1 60.5
-GPT-4o-20240806 / 92.7 87.2 87.6 72.2 30.5 25.0 36.5 61.7
-Open-source Models 7B+ Scale
-[…]
-DeepSeek-V3 / 91.5 86.6 87.6 73.0 36.3 27.7 39.9 63.2
-Llama3.3-70B-Instruct / 84.1 80.5 87.6 73.5 29.1 28.4 28.4 58.8
-Open-source Models $\sim$7B Scale
-[…]
-DSCoder-6.7B-Instruct 2BT 74.4 71.3 74.9 65.6 12.8 10.1 15.5 46.4
-MagiCoder-S-DS-6.7B 185K 76.8 71.3 79.4 69.0 13.4 13.5 12.8 48.0
-OpenCodeInterpreter-DS-6.7B 178K 77.4 72.0 76.5 66.4 7.3 13.5 16.9 47.1
-WaveCoder-Ultra-6.7B 130K 75.0 69.5 74.9 63.5 12.9 12.8 16.9 46.5
-RefineCoder-DS-6.7B   Iter0 80K 73.8 67.7 77.5 65.1 13.6 18.2 10.1 46.6
-RefineCoder-DS-6.7B   Iter1 80K 74.4 68.9 77.0 66.4 14.1 18.9 14.2 47.7+1.1
-RefineCoder-DS-6.7B   Iter2 80K 74.4 70.3 79.6 66.9 14.3 19.6 14.2 48.5+1.9
-RefineCoder-DS-6.7B   Iter3 80K 75.0 70.7 80.2 67.2 14.2 20.3 15.5 49.0+2.4
-[Uncaptioned image]Qwen2.5-Coder-7B-Base / 61.6 53.0 76.9 62.9 / / 16.2      /
-Qwen2.5-Coder-7B-Instruct Millions 88.4 84.1 83.5 71.7 18.1 20.3 20.3 55.2
-RefineCoder-QW-7B   Iter0 20K 85.4 78.0 79.4 65.1 18.2 18.2 20.9 52.2
-RefineCoder-QW-7B   Iter1 20K 86.0 79.3 84.7 71.4 17.4 16.2 18.9 53.4+1.2
-RefineCoder-QW-7B   Iter2 20K 86.6 80.5 84.7 72.0 19.2 18.2 19.6 54.4+2.2
-RefineCoder-QW-7B   Iter3 20K 87.2 81.1 85.2 73.0 19.1 19.6 20.9 55.2+3.0
-Table 1: Main Results.
-LCB/BCB denotes LiveCodeBench/BigCodeBench.
-Inst/Comp denote Instruct/Complete.
-[Uncaptioned image] and [Uncaptioned image] denotes the two base models, and the rest are the instruction models.
-2BT denotes 2B tokens.
+(2024) and BigCodeBench-hard to measure the advanced coding ability of code model.
 4.2 Baselines
-$\circ$ Proprietary Models: Gemini-1.5-Pro-002 Google (2024), Claude-3.5-Sonnet-20240620 Anthropic (2024), and GPT-4o-20240806 OpenAI (2024a).
-$\circ$ Open-source Models 7B+ Scale: Qwen2.5-Coder-32B-Instruct Hui et al. (2024), DeepSeek-V3 DeepSeek-AI (2024), and Llama3.3-70B-Instruct Grattafiori et al. (2024).
-$\circ$ Open-source Models $\sim$7B Scale: DSCoder-6.7B Series Guo et al.
+$\circ$ Based on DSCoder-6.7B-Base: DSCoder-6.7B-Instruct Guo et al.
 (2024), MagiCoder-S-DS-6.7B Wei et al.
@@
 (2024a), WaveCoder-Ultra-6.7B Yu et al.
-(2024), and Qwen2.5-Coder-7B Series Hui et al.
 (2024).
+$\circ$ Based on Qwen2.5-Coder-7B-Base: Qwen2.5-Coder-7B-Instruct Hui et al. (2024).
 4.3 Initial SFT Dataset $D_{0}$
@@
 (2024).
-To ensure the accuracy and reliability of the experimental results, we constructed a high-quality SFT dataset by combining the Evol-Instruct and SelfCodeAlign Wei et al.
+To ensure the accuracy and reliability of the experimental results, we construct a high-quality SFT dataset by combining Evol-Instruct and SelfCodeAlign Wei et al.
 (2024a).
@@
 These code snippets have been preprocessed to ensure high quality and contamination-free.
-Then, we prompt GPT-4o again to iteratively evolve these instructions using Evol-Instruct strategy.
+Then, we prompt GPT-4o again to iteratively evolve these instructions using the Evol-Instruct strategy.
 Finally, we prompt GPT-4o to generate code responses for the instructions.
-Through this pipeline, we obtained an 80K high-quality and diverse python instruction dataset.
-See Appendix C for detailed data construction prompts.
-Dataset Size HumanEval (+) MBPP (+) LCB BCB-hard
-Magic-OSS-Instruct Wei et al. (2024b)  [^1] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K 75K 4.45 9.40 2.63 4.84
-Magic-Evol-Instruct Wei et al. (2024b)  [^2] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K 110K 43.20 19.40 2.91 4.46
-Evol-CodeAlpaca-v1 Luo et al. (2024)  [^3] https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1 110K 47.04 19.46 2.91 4.46
-Code-FeedBack Zheng et al. (2024a)  [^4] https://huggingface.co/datasets/m-a-p/Code-Feedback 68K 30.50 17.67 3.16 4.78
-Ours 80K 4.97 7.00 1.54 5.43
+Through this pipeline, we obtain an 80K high-quality and diverse python instruction dataset.
+See Appendix B for detailed data construction prompts.
+Models Data Size HumanEval MBPP BigCodeBench Average
+Base Plus Base Plus Inst Comp
+[Uncaptioned image] DSCoder-6.7B-Base / 47.6 39.6 72.0 58.7 / 41.8 /
+DSCoder-6.7B-Instruct 2BT 74.4 71.3 74.9 65.6 35.5 43.8 60.9
+MagiCoder-S-DS-6.7B 185K 76.8 71.3 79.4 69.0 36.2 47.6 63.4
+OpenCodeInterpreter-DS-6.7B 178K 77.4 72.0 76.5 66.4 37.1 44.6 62.3
+WaveCoder-Ultra-6.7B 130K 75.0 69.5 74.9 63.5 33.9 43.7 60.1
+RefineCoder-DS-6.7B   Iter0 80K 73.8 67.7 77.5 65.1 37.0 46.6 61.3
+RefineCoder-DS-6.7B   Iter1 80K 74.4 68.9 77.0 66.4 40.6 48.0 62.6↑1.3
+RefineCoder-DS-6.7B   Iter2 80K 74.4 70.3 79.6 66.9 40.6 47.9 63.3↑2.0
+RefineCoder-DS-6.7B   Iter3 80K 75.0 70.7 80.2 67.2 41.1 49.6 64.0↑2.7
+[Uncaptioned image]Qwen2.5-Coder-7B-Base / 61.6 53.0 76.9 62.9    /    /    /
+Qwen2.5-Coder-7B-Instruct Millions 88.4 84.1 83.5 71.7 40.4 48.8 69.5
+RefineCoder-QW-7B   Iter0 80K 84.9 78.1 79.4 65.8 40.0 47.4 65.9
+RefineCoder-QW-7B   Iter1 80K 86.0 79.3 84.1 70.8 41.6 49.0 68.5↑2.6
+RefineCoder-QW-7B   Iter2 80K 86.6 82.5 83.4 71.7 42.3 50.9 69.6↑3.7
+RefineCoder-QW-7B   Iter3 80K 87.2 83.3 85.6 72.9 42.0 51.1 70.4↑4.5
+Table 2: Results on HumanEval, MBPP and BigCodeBench. Values in red indicate severe data leakage.
+Dataset Size HumanEval (+) MBPP (+) LCB BCB
+Magic-OSS-Instruct Wei et al. (2024b) 75K 4.45 9.40 2.63 4.56
+Magic-Evol-Instruct Wei et al. (2024b) 110K 43.20 19.40 2.91 4.10
+Evol-CodeAlpaca-v1 Luo et al. (2024) 110K 47.04 19.46 2.91 4.69
+Code-FeedBack Zheng et al. (2024a) 68K 30.50 17.67 3.16 4.49
+Ours 80K 4.97 7.00 1.54 5.00
 Table 3: The Test Leakage Indicator (TLI) quantifies data leakage by measuring the average maximum n-gram overlap between dataset samples and benchmark samples.
 The larger the TLI value, the more severe the data leakage.
-Values in red indicate severe data leakage.
 4.4 Implement Details
 Iterative Training We employ DSCoder-6.7B-Base and Qwen2.5-Coder-7B-Base as the base pre-trained models.
-To obtain an initial code model $M_{0}$, we fine-tune DSCoder-6.7B-Base on the 80K initial SFT dataset, resulting in RefineCoder-DS-6.7B (Iter0).
-Furthermore, since Qwen2.5-Coder-Base exhibits strong code understanding and generation capabilities, we accelerate our experimental iterations by randomly sampling 20K examples from the same dataset for fine-tune it, yielding RefineCoder-QW-7B (Iter0).
-The two RefineCoder models undergo three iterative training using their respective datasets under our ACR framework.
-In the ACR pipeline, the model self-samples 5 code responses with a temperature of 0.7.
+To obtain an initial code model $M_{0}$, we fine-tune DSCoder-6.7B-Base and Qwen2.5-Coder-Base on the 80K initial SFT dataset, resulting in RefineCoder-DS-6.7B and RefineCoder-QW-7B.
+The two RefineCoder models undergo three iterative training under our ACR framework.
+In the ACR pipeline, the model self-samples 7 code responses with a temperature of 0.7.
 We use Qwen2.5-32B-Instruct Hui et al.
@@
 We set the number of training epochs to 2, the global batch size to 64, and the learning rate to 5e-6, employing the AdamW optimizer with a cosine learning rate decay strategy.
-All our training is performed using 16 A100-80G GPUs, utilizing the LLaMA-Factory Zheng et al.
+All our training uses 16 A100-80G GPUs, utilizing the LLaMA-Factory Zheng et al.
 (2024b).
-Evaluation We use the Pass@1 metric to evaluate the performance of model on the benchmark.
-For baseline results, we prioritize those from the benchmark leaderboard  [^5] https://evalplus.github.io/leaderboard.html  [^6] https://livecodebench.github.io/leaderboard.html  [^7] https://huggingface.co/spaces/bigcode/bigcodebench-leaderboard or the original papers.
+Evaluation We use the Pass@1 metric to evaluate the performance of model on benchmarks.
+We prioritize baseline results from the benchmark leaderboards or the original papers.
 If unavailable, we evaluate them locally using the same settings as the RefineCoder series.
 4.5 Experimental Results
-As shown in Table 1, the RefineCoder series achieves impressive performance gains on various code generation benchmarks.
-While the initial RefineCoder-DS-6.7B and RefineCoder-QW-7B exhibit moderate performance, the two models achieve an average Pass@1 improvement of 2.4 and 3.0, respectively, on all benchmarks after three iterations of refinement.
-Furthermore, it is worth noting that the RefineCoder series is fine-tuned with a very limited amount of data (80K or 20K).
-Despite this, the RefineCoder models (Iter3) still outperform or match baselines of the same size, which are fine-tuned with more extensive data.
-Concretely, data refinement improves the fundamental programming abilities of the RefineCoder series, as demonstrated by continuous performance gains over iterations on HumanEval, HumanEval+, and MBPP+.
-Notably, RefineCoder-QW-7B shows a remarkable 7.9 point increase (Iter0 -> Iter3) in Pass@1 on MBPP+, highlighting the efficacy of this approach.
-In the more challenging benchmarks, LiveCodeBench and BigCodeBench-hard, the performance of the RefineCoder also improves after iterations.
-Compared to the baseline OpenCodeInterpreter, RefineCoder-DS-6.7B (Iter2) outperforms it by 7 points on LiveCodeBench, illustrating the strong competitive programming abilities of the RefineCoder series.
-These results demonstrate that the model can iteratively enhance code generation capabilities by self-generated code and critique with the help of our ACR method.
+As shown in Table 1, the RefineCoder series achieves impressive performance gains on two challenging code benchmarks.
+While the initial RefineCoder-DS-6.7B and RefineCoder-QW-7B exhibit moderate performance, the two models achieve an average Pass@1 improvement of 2.7 and 3.1, respectively, after three refinement iterations.
+Notably, the performance improved by 2.0 and 2.3 points from iter0 to iter1, highlighting that incorporating self-generated code and external critique yields better results than directly distilling code responses from the teacher model.
+Compared to baselines of the same size, the RefineCoder-DS-6.7B and RefineCoder-QW-7B (Iter3) outperform them by 3.6 and 2.3 points, respectively, using only 80K data.
+This demonstrates the superiority of our method in improving advanced coding ability.
+Similarity, our method also improves the basic coding ability, as shown in Table 2.
+After three iterations, our model achieves average performance improvements of 2.7 and 4.5 points, respectively, on HumanEval(+), MBPP(+) and BigCodeBench.
+This demonstrates that the model is capable of identifying and addressing deficiencies by learning from both its self-generated code and external feedback.
+Compared with all baselines, our models achieve the best average Pass@1 performance.
+Nevertheless, we find that the performance of RefineCoders on the Humaneval(+) are inferior to the state-of-the-art baseline.
+This discrepancy prompts us to conduct a data leakage analysis of the datasets in the next section.
 5 Further Study
 5.1 Data Leakage Analysis
-In Table 1, we observe that while the RefineCoder (Iter3) surpasses or matches the baselines on average pass@1, it still lags behind the state-of-the-art baseline of the same size on HumanEval (+) and MBPP+.
+In Table 2, we observe that while the RefineCoder (Iter3) surpasses the baselines on average pass@1, it still lags behind the state-of-the-art baseline on HumanEval(+) and MBPP+.
 To better explore this performance gap, we investigate potential data leakage in the datasets used by the baselines.
-Concretely, we collect the open-source datasets used by MagiCoder-S-DS-6.7B, OpenCodeInterpreter-DS-6.7B, and WaveCoder-Ultra-6.7B (the datasets for the other baselines are not public) and evaluate data leakage using the TLI metric proposed by Wang et al.
+Concretely, we analyze the four open-source datasets  [^1] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K  [^2] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K  [^3] https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1  [^4] https://huggingface.co/datasets/m-a-p/Code-Feedback with the help of the TLI metric proposed by Wang et al.
 (2024b).
+These datasets are used by MagiCoder-S-DS-6.7B, OpenCodeInterpreter-DS-6.7B, and WaveCoder-Ultra-6.7B (The datasets for DSCoder-6.7B-Instruct and Qwen2.5-Coder-7B-Instruct are not publicly available).
 The results are shown in Table 3, Magic-OSS-Instruct and our dataset are entirely free from data leakage.
-In contrast, Magic-Evol-Instruct, Evol-CodeAlpaca-v1, and Code-FeedBack demonstrate severe data leakage with respect to HumanEval (+) and MBPP (+).
-These datasets are used by the three aforementioned baselines, making it challenging for RefineCoder to surpass these baselines.
-We also present a intuitive scatter plot of similarity scores in the appendix E for a more detailed analysis.
+In contrast, Magic-Evol-Instruct, Evol-CodeAlpaca-v1, and Code-FeedBack demonstrate severe data leakage concerning HumanEval (+) and MBPP (+).
+These three datasets are used by three baselines, making it challenging for RefineCoder to surpass them.
+We also present an intuitive scatter plot of similarity scores in the appendix D for a more detailed analysis.
 HumanEval+ (%) Best Worst Error Rate
 RefineCoder-DS-6.7B Iter3 76.2 60.4-15.8 4.3
-RefineCoder-QW-7B Iter3 81.7 75.0-6.7 3.0
+RefineCoder-QW-7B Iter3 83.6 75.0-8.6 3.1
 MBPP+ (%) Best Worst Error Rate
-RefineCoder-DS-6.7B Iter3 65.3 60.1-5.2 6.3
-RefineCoder-QW-7B Iter3 70.4 69.6-0.8 5.8
+RefineCoder-DS-6.7B Iter3 70.3 60.1-10.2 5.8
+RefineCoder-QW-7B Iter3 73.4 71.1-2.3 5.3
 Table 4: Best and Worst denote the performance of code with the highest and lowest scores.
-Error rate denotes the proportion of cases where the best code fails the test cases, but the worst code passes.
+Error rate denotes the proportion of cases where the best code fails the test cases but the worst code passes.
 5.2 The effectiveness of Composite Scoring System
-The performance improvement through iterative training preliminarily validates the effectiveness of the scoring system.
-In this section, we further verify it by enabling the model sampling 10 code responses for each programming question in benchmarks and scoring them.
+The performance improvement through iterative training preliminarily validates the effectiveness of our scoring system.
+In this section, we further verify it by enabling the model to sample 10 code responses for each programming question in benchmarks and scoring them.
 As shown in Table 4, the performance of the highest-scoring code far exceeds that of the lowest-scoring code, while the scoring system maintains a low error rate.
 This clearly validates its effectiveness.
-5.3 The effectiveness of Selective Critique Strategy
-To validate the effectiveness of the selective critique strategy in ACR, we conduct ablation experiments.
-As shown in Figure 3, for each iteration, we construct two variants of the RefineCoder without the 3.1 module in Figure 2, and another without the 3.2 module.
+Figure 3: Ablation Study. The y-axis denotes the average pass@1 value on LiveCodeBench and BigCodeBench-hard.
+Figure 4: Evaluation with feedback using RefineCoder-QW-7B.
+5.3 The effectiveness of Selective Critiquing Strategy
+We conduct ablation experiments to validate the effectiveness of the selective critiquing strategy in ACR.
+As shown in Figure 3, for each iteration, we construct two variants of RefineCoder (without 3.1 module of Figure 2 or without 3.2 module).
 During the iteration process, ablating either module leads to a performance drop in both RefineCoder-QW-7B and RefineCoder-DS-6.7B, indicating the effectiveness of our strategy.
-Furthermore, removing the second-round critique data has a greater impact on performance.
-We believe this action prevents the model from reflecting on self-generated low-quality codes and external critiques, thereby hindering performance improvement.
-Figure 3: Ablation Study. The y-axis denotes the average pass@1 value on all benchmarks.
+Furthermore, removing the second-turn critique data has a greater impact on performance.
+We believe this action prevents the model from reflecting on self-generated low-quality code and external critique, thereby hindering performance improvement.
 5.4 Multilingual Code Generation
-Figure 4: Evaluation with feedback using RefineCoder-QW-7B.
 Models C++ C# Java Bash TypeScript JavaScript AVG
@@
 RefineCoder-DS-6.7B   Iter3 63.4 58.2 57.0 39.2 62.9 65.8 57.8
-RefineCoder-QW-7B   Iter0 60.2 73.4 68.4 49.4 78.0 78.3 68.7
-RefineCoder-QW-7B   Iter1 62.7 74.1 65.2 51.3 76.1 77.0 68.6
-RefineCoder-QW-7B   Iter2 63.8 74.1 69.6 48.7 78.0 77.6 69.4
-RefineCoder-QW-7B   Iter3 60.2 75.3 73.4 53.2 79.2 77.6 70.6
+RefineCoder-QW-7B   Iter0 60.7 72.4 68.0 49.9 77.6 78.1 67.8
+RefineCoder-QW-7B   Iter1 62.1 74.6 65.0 50.8 76.8 77.4 67.8
+RefineCoder-QW-7B   Iter2 63.7 74.1 69.6 49.7 78.3 77.7 68.9
+RefineCoder-QW-7B   Iter3 60.9 75.3 73.8 52.2 79.3 77.9 69.9
 Table 5: Performance of RefineCoder on the MultiPL-E.
-As shown in Table 4, we also evaluate the out-of-distribution (OOD) code generation performance on multilingual benchmark MultiPL-E Cassano et al.
-(2022), despite fine-tuning on a Python-only dataset.
-After one and three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve optimal average performance, surpassing the initial model by 1.0 and 1.9 points, respectively.
-In particular, RefineCoder-QW-7B (Iter3) achieves best results among half of the programming languages.
+As shown in Table 5, we also evaluate the out-of-distribution (OOD) code generation performance on the multilingual benchmark MultiPL-E Cassano et al.
+(2022) despite fine-tuning on a Python-only dataset.
+After one and three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve optimal average performance, surpassing the initial model by 1.0 and 2.1 points, respectively.
+In particular, RefineCoder-QW-7B (Iter3) achieves the best results among half of the programming languages.
 This demonstrates that the effectiveness of our ACR method generalizes well to the OOD code generation domain.
@@
 Following  Zheng et al.
-(2024a), we design two types of external feedback to evaluate this ability of RefineCoder: 1) Execution Feedback: Model refines its self-generated erroneous code based on execution results from the executor.
-2) Human Feedback: GPT-4o first analyzes the programming question, initial error code, and execution results to generate improvement suggestions that mimic human thinking.
-Model then refines the code based on these suggestions.
+(2024a), we design two types of external feedback to evaluate this ability of RefineCoder: 1) Execution Feedback: Model refines its self-generated erroneous code based on execution results from the executor; 2) Human-like Feedback: GPT-4o first analyzes the programming question, initial error code, and execution results to generate improvement suggestions that mimic human thinking.
+The model then refines the code based on these suggestions.
 The results of RefineCoder-QW-7B on HumanEval(+) and MBPP(+) are shown in Figure 4.
-It can be observed that external feedback improves performance on all benchmarks, with human feedback yielding a more significant enhancement.
+It can be observed that external feedback improves performance on all benchmarks, with human-like feedback yielding a more significant enhancement.
 Surprisingly, with the help of human-like improvement suggestions, RefineCoder-QW-7B (Iter1 and Iter3) outperforms GPT-4o on three benchmarks, demonstrating its ability to understand feedback and refine code.
-Models HumanEval MBPP
-Base Plus Base Plus
-Qwen2.5-Coder-7B-Base 61.6 53.0 76.9 62.9
-     +DPO 59.8 52.4 44.2 32.6
-     +ORPO 85.4 79.9 84.4 70.4
-RefineCoder-QW-7B Iter0 85.4 78.0 79.4 65.1
-     +DPO 61.0 50.0 60.6 42.3
-     +ORPO 86.6 79.3 81.7 69.3
-Table 5: Performance comparison of different preference optimization strategies in reinforcement learning.
-5.6 Preference Optimization
-During the iterative process, a substantial amount of preference data is generated.
-For each programming instruction, we select the top-1 and bottom-1 ranked responses to form preference pairs, thereby exploring the potential of preference learning.
-As shown in Table 5, we use DPO Rafailov et al.
-(2023) and ORPO Hong et al.
-(2024) to optimize the base and SFT models.
-We observe the opposite results due to the differences in preference strategies.
-DPO leads to a significant performance decline due to the intrinsic risk of reducing the likelihood of correct code during training Feng et al.
-(2024); Tao et al.
-(2024); Pal et al.
-(2024).
-In contrast, ORPO builds upon the SFT loss to maximize the likelihood of correct code, resulting in improved performance.
-The results are consistent with those presented in the CodeLutra Tao et al.
-(2024), reflecting the indispensability of SFT loss and the critical impact of preference strategy design on performance.
 6 Conclusions
@@
 We design a composite scoring system and a selective critiquing strategy.
-These two components are centred around LLM-as-a-Judge and LLM-as-a-Critic to evaluate and critique self-generated code.
+These two components are centered around LLM-as-a-Judge and LLM-as-a-Critic to evaluate and critique self-generated code.
 This simulates the process where a student solves a problem independently and then refines it by comparing it with the reference answer.
-We develop the RefineCoder-DS-6.7B and RefineCoder-QW-7B models and demonstrate the effectiveness of iterative ACR on HumanEval (+), MBPP (+), LiveCodeBench, and BigCodeBench-hard.
-Further studies reveal the impact of the components in ACR and demonstrate its OOD code generation ability.
+We develop the RefineCoder-DS-6.7B and RefineCoder-QW-7B models and demonstrate the effectiveness of iterative ACR on multiple code generation benchmarks.
+Further studies reveal the impact of the each component in ACR.
 7 Limitations
@@
 Furthermore, while ACR can apply to other reasoning-intensive tasks, such as mathematics, this paper has not fully explored these domains.
-Appendix A Elo Algorithm
-Firstly, we initialize Elo score $E^{i}=1000$ for each code response $y^{i}$.
-Then, we iteratively update the Elo scores by using the relative scores between any two responses.
-Taking the response pair $y^{i}$ and $y^{j}$ as an example, we obtain their match results $S^{i}$ and $S^{j}$:
-$S^{i}=\frac{\textrm{score}_{judge}^{ij}}{\textrm{score}_{judge}^{ij}+\textrm{score}_{judge}^{ji}},$ (8)
-$S^{j}=\frac{\textrm{score}_{judge}^{ji}}{\textrm{score}_{judge}^{ij}+\textrm{score}_{judge}^{ji}}.$ (9)
-Then we obtain the expected probabilities of winning $P^{ij}$ and $P^{ji}$ for $y^{i}$ and $y^{j}$:
-$P^{ij}=\frac{1}{1+10^{\frac{E^{j}-E^{i}}{400}}},$ (10)
-$P^{ji}=1-P^{ij}.$ (11)
-Finally, we update the Elo score:
-$E^{i}_{new}=E^{i}+K\times(S^{i}-P^{ij}),$ (12)
-$E^{j}_{new}=E^{j}+K\times(S^{j}-P^{ji}).$ (13)
-where $K=32$ is a adjustment factor. After the iteration, we obtain the Elo score for each code response and then normalize it to derive the final judge score:
-$\textrm{score}^{i}_{judge}=\frac{E_{i}-\min(E)}{\max(E)-\min(E)},$ (14)
-where $i\in[0,n]$ and $\textrm{score}^{i}_{judge}\in[0,1]$.
-Appendix B Prompts for Judge and Critic
-The prompt for pairwise LLM-as-a-Judge and LLM-as-a-Critic are shown in Figure 7 and  8.
-Appendix C Prompts for constructing SFT Dataset
-We called GPT-4o three times to create the SFT dataset, with the following prompts used:
+Appendix A Prompts for Judge and Critic
+The prompt for point/pair-wise LLM-as-a-Judge and LLM-as-a-Critic are shown in Figure 8,  9 and  10.
+Appendix B Prompts for Constructing SFT Dataset
+We called GPT-4o to create the SFT dataset, with the following prompts used:
 $\circ$ Prompt for generating code-related concepts and programming instructions, as shown in Figure 11.
-$\circ$ Prompts for evolving existing instructions, as shown in Figures 10,  11,  1213141516,  17,  18.
+$\circ$ Prompts for evolving existing instructions, as shown in Figures 12,  13,  1415161718,  19,  20.
 $\circ$ For generating code responses, We directly feed the instructions to the model.
-Appendix D Prompts for Evaluation
+Appendix C Prompts for Evaluation
 For Humaneval (+) and MBPP (+), we use the prompts designed by OpenCodeInterpreter Zheng et al.
@@
 When conducting evaluations with execution feedback, we have designed a prompt as illustrated in Figure 21.
-When executing human feedback, following OpenCodeInterpreter, we generate two prompts: one for generating improvement suggestions and the other for refine code according to human feedback, as shown in Figures 21 and  20.
-Appendix E Data Leakage Analysis
+When executing human-like feedback, following OpenCodeInterpreter, we generate two prompts: one for generating improvement suggestions and the other for refine code according to human-like feedback, as shown in Figures 22 and  23.
+Appendix D Data Leakage Analysis
 We use 4-grams and 5-grams to compute the similarity between the instructions in the training set and the questions in the benchmark.
@@
 In contrast, on the leakage-free LiveCodeBench and BigCodeBench-hard benchmarks, our model can outperform all baselines.
-Appendix F Scaling Law of Iterative ACR
-We study the performance trends of RefineCoder-DS-6.7B obtained using ACR and the standard SFT method as the dataset size increases.
-For the SFT method, we directly trained the model using four datasets of different sizes.
-For RefineCoder series, we employ an iterative training strategy.
-For example, after obtaining RefineCoder-DS-10k by fing-tuning on the 10k dataset, we apply ACR to update this 10k dataset, and then randomly sample another 10k from the data pool to form a 20k dataset.
-The model is fine-tuned again to obtain RefineCoder-DS-20k.
-The experimental results are shown in Figure 6, as the dataset size increases, our model achieves greater performance improvements compared to the SFT model on HumanEval and MBPP.
-Moreover, the performance gap between the two models does not show any signs of narrowing, underscoring the impressive scaling capabilities of the iterative ACR method.
+Appendix E Diminishing Returns
+As shown in Figure 5, we present the proportions of two types of data during the iterative process of RefineCoder-QW-7B.
+The self-generated code struggles more to outperform as iterations continue, indicating diminishing potential for further gains through self-refinement.
+The corresponding experimental results are shown in Figure 6.
+As observed, the model exhibits clear diminishing returns after the third iteration.
+This suggests that the model may have entered a mild overfitting stage, where further refinement not only fails to improve performance but may also exacerbate the issue.
+Figure 5: Proportions of two types of data in the iterative process.
+Figure 6: The average Pass@1 value of RefineCoder-QW-7B on HumanEval(+), MBPP(+) and BigCodeBench.
 Figure 7: Scatter plot of similarity scores between the datasets and four benchmarks, with average similarity scores for different datasets indicated by horizontal dashed lines.
@@
 Different y-axis ranges are set for better visualization.
-Figure 6: Performance curves of RefineCoder obtained using iterative ACR and the standard SFT model as the dataset size increases.
-Figure 7: The prompt template for pairwise LLM-as-a-Judge.
-Figure 8: The prompt template for pairwise LLM-as-a-Critic.
+Figure 8: The prompt template for pointwise judgment.
+Figure 9: The prompt template for pairwise judgment.
+Figure 10: The prompt template for pairwise critique.
 Figure 11: The prompt template for generating code-related concepts and instructions.
@@
 Figure 21: The prompt template for Execution Feedback.
-Figure 20: The prompt template for Human Feedback.
 Figure 22: The prompt template for generating improvement suggestions.
+Figure 23: The prompt template for Human-like Feedback.